OpenStudy (anonymous):

solve word problem of quadratic equation in acertain positive fraction the denominator is greater than the numerator by 3.if 1 is subtracted from the fraction it reduces by 1/14 find the fraciton

OpenStudy (anonymous):

let the num be x and denom be y then original fraction is x/y according to question y = x + 3 x so original fraction is = ------- x + 3 on subtracting 1 frm fraction it reduces by 1/14 x x x so ------- - 1 = ------ - ---------- x + 3 x + 3 14(x + 3) x -1(x + 3) x(14) - x ---------- = ----------- x + 3 14(x + 3) x - x - 3 14x -x -------- = --------- x + 3 14x +42 -3 13x ------- = ---------- x + 3 14x + 42

OpenStudy (anonymous):

cross multiplying -3(14x + 42) = 13x (x + 3) -42x - 126 = 13x^2 + 39x 13x^2 + 39x + 42x + 126 = 0 13x^2 + 81x + 126 = 0 solve this and you will get answer

OpenStudy (anonymous):

@tsigma the question says "reduces by 1/14" and not reduces to 1/14 so I guess yr working and solutions are not correct.......☺

OpenStudy (anonymous):

Hoever if the person sending the question has made a mistake, then your solution is ok....

OpenStudy (anonymous):

" -3(14x + 42) = 13x (x + 3)" Ditch x+3 both sides.